OpenStudy (anonymous):

solve for x and y 2x − y = −5 −5x + y = 2 help please ?

OpenStudy (bahrom7893):

from the top equation: 2x + 5 = y plug that into the bottom: -5x + 2x + 5 = 2 -3x = -3 x = 1

OpenStudy (bahrom7893):

y = 7

OpenStudy (anonymous):

I don't know how you went from 2x − y = −5 to 2x + 5 = y. I'm guessing that the asker might not either